OpenStudy (anonymous):

how do i find the derivative of x^2-2x-3/x+2

OpenStudy (anonymous):

for the first part (x^2) you just use the power rule

OpenStudy (anonymous):

for the second part you use (-2x-3/x+2) you use the quotient rule.

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x^2-2x-3,f'(x)+2x-2,g(x)=x+2,g'(x)=1\]

OpenStudy (anonymous):

sorry i meant \(f'(x)=2x-2\)